OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle cos 107° cos 42° + sin 107° sin 42° sin 65° cos 65° cos 149° sin 149°

OpenStudy (anonymous):

Use cosA cosB + sinAsinB = cos(A-B)

OpenStudy (anonymous):

can you give me the answer please

OpenStudy (anonymous):

cos149

OpenStudy (anonymous):

thank you

OpenStudy (raden):

that is A+B, not A-B @sachinrajsharma

OpenStudy (raden):

cos(107-42) = cos ?

OpenStudy (raden):

107-42 = ...

OpenStudy (anonymous):

B. cos 65